# ZK-ACE: Identity-Centric Zero-Knowledge Authorization for Post-Quantum Blockchain Systems

### 09 Mar 2026 v1

@@ 1.5 Contributions
    We provide a structural, protocol-level accounting of on-chain authorization data requirements, demonstrating an order-of-magnitude reduction relative to signature-based post-quantum authorization, without relying on experimental benchmarks.
@@ 1.7 Paper Organization
Section 2 surveys related work. Section 3 introduces notation, the assumed DIDP interface, and the threat model. Section 4 describes the system model and transaction flow. Section 5 defines the identity commitment and binding rules. Section 6 clarifies the authorization semantics. Section 7 specifies the ZK circuit. Section 8 presents formal security definitions and proofs. Section 9 discusses on-chain verification and deployment. Section 10 provides the structural data accounting. Section 11 describes scalability extensions. Section 12 addresses limitations, and Section 13 concludes.
@@ 2 Related Work
STARKs [9] and FRI-based interactive oracle proofs provide transparent, plausibly post-quantum-secure proof generation without a trusted setup. Recent work by Chiesa, Fenzi, and Weissenberg [25] constructs zero-knowledge IOPPs for constrained interleaved codes with near-optimal $(1+o(1))$ soundness-to-proof-length overhead, further advancing the efficiency of hash-based proof systems. These developments improve the proof system layer and are complementary to ZK-ACE, which operates at the authorization model layer. Even with an optimal transparent proof system, embedding lattice-based signature verification into a ZK circuit still requires expressing high-dimensional polynomial arithmetic (NTTs, non-native modular reductions) as arithmetic constraints—yielding circuits on the order of millions of constraints. ZK-ACE sidesteps this bottleneck entirely: its circuit requires only a small number of ZK-friendly hash evaluations ($\approx$}\,1,100–1,400 R1CS constraints in the base configuration, up to $\approx$}\,1,800 with additional multi-element encoding; see Section 7.3), and is compatible with any backend proof system, including hash-based and transparent constructions.
@@ 7.3 Core Constraints
The circuit cost is dominated by the in-circuit hash evaluations. Table 1 provides a representative breakdown using Poseidon with a standard parameterization (width $t=3$, 8 full rounds over the BN254 scalar field), where each hash invocation requires approximately 250–300 R1CS constraints [6].
Table 1: Representative constraint breakdown for ZK-ACE circuit (Poseidon, $t=3$, BN254).
Constraint Hash calls Est. R1CS
(C1) Commitment consistency 1 250–300
(C2) Derivation correctness 1 250–300
(C3) Authorization binding 1 250–300
(C4) Replay prevention 1 250–300
Aux. equality + range checks  ${\approx}\,100$–200
Total 4 $\approx$}\,1,100–1,400
Domain separation (C5) is absorbed into the inputs of the preceding hashes and does not incur additional hash invocations. Depending on the number of input field elements per Poseidon call (which affects padding and multi-round absorption), actual constraint counts may vary; an additional Poseidon invocation for multi-element encoding would bring the total to approximately 1,400–1,800 R1CS constraints.
For comparison, verifying a lattice-based signature such as ML-DSA inside a zero-knowledge circuit requires expressing: (i) Number Theoretic Transforms (NTTs) over degree-256 polynomial rings in $\mathbb{Z}_{q}$ with $q=2^{23}-2^{13}+1$, each requiring $O(n\log n)$ multiplications in a non-native field; (ii) high-bit-width modular reductions (23-bit modulus emulated over a $\approx$}\,254-bit proof-system field, requiring range-check gadgets per limb); and (iii) rejection-sampling and hint-reconstruction logic with data-dependent branching. ML-DSA verification at NIST level 2 involves $k\times\ell=4\times 4$ matrix–vector NTTs and $k=4$ polynomial reconstructions, yielding a structural lower bound on the order of millions of R1CS constraints even with optimized non-native arithmetic. Based on this structural analysis, the ZK-ACE circuit is estimated to be roughly three orders of magnitude smaller than an in-circuit post-quantum signature verification. This estimate reflects the structural gap between a small number of ZK-friendly hash invocations and the full algebraic machinery of lattice-based signature verification; precise ratios will depend on the circuit toolchain and optimization level used in a concrete implementation. The resulting circuit size makes proving feasible on consumer-grade hardware and enables practical batch and recursive composition (Section 11).
@@ 12 Limitations and Future Work
  • Foundations-level treatment. This work establishes the theoretical framework—formal definitions, security proofs, and structural data accounting—rather than providing a concrete circuit implementation or gas-precise, chain-specific benchmarks. The R1CS constraint estimates (Section 7.3) are derived from structural analysis of the required Poseidon invocations and should be validated against a concrete circuit implementation (e.g., Circom, gnark) for any specific deployment. Such implementation-level evaluation is deferred to follow-up work.
  • Prover cost. While verifier cost is reduced, prover cost (ZK proof generation) remains non-trivial, particularly on resource-constrained devices. Hardware acceleration and circuit optimization are active areas of research.
@@ 13 Conclusion
The structural data accounting shows an order-of-magnitude reduction in per-transaction authorization data relative to direct post-quantum signature deployment—achieved not by compressing signatures, but by eliminating the need for signature objects in the consensus path entirely. With support for batch aggregation and recursive composition, ZK-ACE provides a scalable authorization architecture for post-quantum blockchain systems.

### 13 Mar 2026 v2

@@ 1.5 Contributions
    We provide a structural, protocol-level accounting of on-chain authorization data requirements, demonstrating an order-of-magnitude reduction relative to signature-based post-quantum authorization.
  8. 8.
    We report a reference implementation and single-threaded CPU benchmarks validating the constraint counts and demonstrating sub-millisecond verification and sub-100 ms proving on consumer hardware.
@@ 1.7 Paper Organization
Section 2 surveys related work. Section 3 introduces notation, the assumed DIDP interface, and the threat model. Section 4 describes the system model and transaction flow. Section 5 defines the identity commitment and binding rules. Section 6 clarifies the authorization semantics. Section 7 specifies the ZK circuit. Section 8 presents formal security definitions and proofs. Section 9 discusses on-chain verification and deployment. Section 10 provides the structural data accounting. Section 11 describes scalability extensions. Section 12 reports a reference implementation and benchmarks. Section 13 addresses limitations, and Section 14 concludes.
@@ 2 Related Work
STARKs [9] and FRI-based interactive oracle proofs provide transparent, plausibly post-quantum-secure proof generation without a trusted setup. Recent work by Chiesa, Fenzi, and Weissenberg [25] constructs zero-knowledge IOPPs for constrained interleaved codes with near-optimal $(1+o(1))$ soundness-to-proof-length overhead, further advancing the efficiency of hash-based proof systems. These developments improve the proof system layer and are complementary to ZK-ACE, which operates at the authorization model layer. Even with an optimal transparent proof system, embedding lattice-based signature verification into a ZK circuit still requires expressing high-dimensional polynomial arithmetic (NTTs, non-native modular reductions) as arithmetic constraints—yielding circuits on the order of millions of constraints. ZK-ACE sidesteps this bottleneck entirely: its circuit requires only a small number of ZK-friendly hash evaluations ($\approx$}\,4,000 R1CS constraints in the reference Poseidon parameterization; see Section 7.3), and is compatible with any backend proof system, including hash-based and transparent constructions.
@@ 7.3 Core Constraints
The circuit cost is dominated by the in-circuit hash evaluations. Table 1 provides the measured constraint breakdown from the reference implementation using Poseidon with width $t=3$, S-box exponent $\alpha=17$, 8 full rounds, and 57 partial rounds over the BN254 scalar field. [^2] The original version of this paper cited the Poseidon paper’s [6] theoretical estimate of  ${\approx}\,250$300300 constraints per hash invocation, which assumes $\alpha=5$ and fewer partial rounds. The reference implementation uses $\alpha=17$ with 57 partial rounds for the BN254 scalar field.
The per-invocation cost scales with input arity: the arkworks sponge gadget (rate r=2r\!=\!2) requires $\lceil k/r\rceil$ absorption rounds for $k$ field-element inputs, at  ${\approx}\,400$ R1CS constraints per round. (C2) requires two hash invocations—an inner hash over 4 inputs plus an outer hash over 1 input. (C3) absorbs 7 field elements ($\mathsf{REV}\|\textsf{AlgID}\|\textsf{Domain}\|\textsf{Index}\|\mathsf{TxHash}\|\mathsf{domain}\|\mathsf{nonce}$), requiring 4 absorption rounds and accounting for 40% of the total circuit.
Table 1: Measured constraint breakdown for the ZK-ACE circuit (Poseidon, $t=3$, $\alpha=17$, BN254, arkworks).
Constraint Inputs Hash calls R1CS
(C1) Commitment consistency 3 1 805
(C2) Derivation correctness 4 + 1 2 1,200
(C3) Authorization binding 7 1 1,615
(C4) Replay prevention 2 1 400
(C5) Domain sep. + enforce_equal 4
Total 5 4,024
Both replay modes (nonce-registry and nullifier-set) produce identical constraint counts. Domain separation (C5) is enforced as an equality constraint on the domain component of $\mathsf{Ctx}$ and adds negligible overhead.
For comparison, verifying a lattice-based signature such as ML-DSA inside a zero-knowledge circuit requires expressing: (i) Number Theoretic Transforms (NTTs) over degree-256 polynomial rings in $\mathbb{Z}_{q}$ with $q=2^{23}-2^{13}+1$, each requiring $O(n\log n)$ multiplications in a non-native field; (ii) high-bit-width modular reductions (23-bit modulus emulated over a $\approx$}\,254-bit proof-system field, requiring range-check gadgets per limb); and (iii) rejection-sampling and hint-reconstruction logic with data-dependent branching. ML-DSA verification at NIST level 2 involves $k\times\ell=4\times 4$ matrix–vector NTTs and $k=4$ polynomial reconstructions, yielding a structural lower bound on the order of millions of R1CS constraints even with optimized non-native arithmetic. The measured ZK-ACE circuit (4,024 R1CS constraints; Table 1) is therefore roughly three orders of magnitude smaller than an in-circuit post-quantum signature verification. This gap reflects the structural difference between a small number of ZK-friendly hash invocations and the full algebraic machinery of lattice-based signature verification. The resulting circuit size makes proving feasible on consumer-grade hardware ($\approx$}\,63 ms single-threaded; see Section 12) and enables practical batch and recursive composition (Section 11).
@@ 11.2 Recursive Proof Composition
12 Reference Implementation and Benchmarks
To validate the structural analysis of Sections 7.3 and 10, we provide a reference implementation of the ZK-ACE circuit and report single-threaded CPU benchmarks. The implementation is open-source and uses the arkworks library ecosystem [26] for BN254 arithmetic and Groth16 proving.
12.1 Implementation
The reference circuit implements all five constraints (C1)–(C5) using the arkworks PoseidonSpongeVar gadget with width $t=3$, S-box exponent $\alpha=17$, 8 full rounds, and 57 partial rounds over the BN254 scalar field. Groth16 [8] is used as the proof system. The circuit produces identical constraint counts (4,024 R1CS) for both the nonce-registry and nullifier-set replay modes, confirming that the replay-mode selection does not affect circuit complexity.
The implementation additionally includes a native (off-chain) Poseidon hash using identical parameters, ensuring functional equivalence between the in-circuit computation and the off-chain witness generation.
12.2 Benchmarks
All measurements were collected using Criterion.rs on an Apple M3 Pro (12-core, 36 GB RAM) running macOS 15.7, Rust 1.92, with single-threaded execution and default compiler optimizations (--release).
ZK proof operations.
Table 4 reports the core Groth16 operations for a single authorization proof. All timings are Criterion.rs medians from cargo bench -p zk-ace --bench prove_verify_bench.
Table 4: Groth16 proof-system benchmarks for ZK-ACE (single-threaded, Apple M3 Pro).
Operation Median 95% CI
Trusted setup (one-time) 120 ms [111, 138] ms
Prove (per transaction) 63 ms [61, 65] ms
[…]
Attestation and pipeline operations.
Table 5 reports timings for the identity-layer primitives and the consensus-critical pipeline stages, measured on the ACE runtime using cargo bench -p ace-runtime --features test-utils --bench pipeline_bench. Note: Phase 2 (proof generation) uses a deterministic mock prover that simulates per-transaction SHA-256 hashing and tree aggregation; it does not invoke real Groth16 proving. The mock-prover timings reflect the overhead of the proving orchestration—witness marshalling, tree construction, and data serialization—exclusive of the ZK proof computation itself, which is benchmarked separately in Table 4.
Table 5: ACE runtime pipeline benchmarks (single-threaded, Apple M3 Pro). Phase 2 uses a mock prover; real Groth16 timings are in Table 4.
Operation Batch size Median time
[…]
Identity commitment ($\mathsf{ID_{com}}$) 1 270 ns
[…]
Block build 2,000 1.97 ms
Phase 2: mock prove 2,000 4.16 ms
End-to-end pipeline 2,000 7.56 ms
REV Argon2id seal/unseal 1 3.5 ms
Mock prover: orchestration overhead only, not real ZK proof generation.
At batch size 2,000, the consensus-critical Attest–Execute path (Phase 1a + 1b + Merkle root) completes in $\approx$}\,0.92 ms, well under 1% of a 400 ms slot window. The attestation-check phase processes each transaction in $\approx$}\,124 ns, consistent with the sub-microsecond HMAC-SHA256 verification expected from the design. The end-to-end pipeline (with mock proving) takes $\approx$}\,7.56 ms; replacing the mock prover with real Groth16 proving would add $\approx$}\,63 ms $\times$ 2,000 of prover work, which is parallelizable across GPU cores and off the critical consensus path (Section 11.1).
@@ 13 Limitations and Future Work
  • Single-platform benchmarks. The benchmarks in Section 12 are collected on a single platform (Apple M3 Pro, single-threaded CPU). Gas-precise, chain-specific cost estimates and GPU-accelerated proving times remain to be measured. Cross-platform validation (e.g., x86, ARM server, WASM) and formal circuit audits are deferred to follow-up work.
[…]
@@ 14 Conclusion
The structural data accounting shows an order-of-magnitude reduction in per-transaction authorization data relative to direct post-quantum signature deployment—achieved not by compressing signatures, but by eliminating the need for signature objects in the consensus path entirely. A reference implementation confirms a total circuit size of 4,024 R1CS constraints with sub-millisecond verification and sub-100 ms single-threaded proving on consumer hardware. With support for batch aggregation and recursive composition, ZK-ACE provides a scalable authorization architecture for post-quantum blockchain systems.